Week of 060605

Open Actions from last week:
  • Install new vomscert on LFC/FTS (SMOD) - DONE

Chair: James

Monday:

Log: Nothing - CERN Holiday

New Actions:

Discussion:

Tuesday:

Log: LFC_SLOWREADDIR over weekend

New Actions:

  • Look into LFC SLOWREADDIR (Jamesc) - DONE
  • Investigate ce102 VM_KILLs (Thorsten) -DONE

Discussion:

Wednesday

Log: RAS

Actions:

  • Add trust rules on LFC nodes for LHCb (Diana/David) -DONE
  • Ops VO needs to be added (Thorsten/Harry) - DONE

Discussion:

  • ATLAS seem to be doing not very much. Harry to check on progress

Thursday

Log: nothing

Actions: * Send OPS YAIM snippet to Thorsten (James) - DONE

Discussion:

Friday

Log: CE101 now has GRIS_WRONG alarms too. Alarm on GRIDVOMS during scheduled intervention - machine should be put into maintenance before carrying out intervention.

Actions: * Add OPS VO to FTS (Gavin) Discussion:

Edit | Attach | Watch | Print version | History: r7 < r6 < r5 < r4 < r3 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r7 - 2007-02-19 - FlaviaDonno
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback